OpenStudy (lgbasallote):

\(\LARGE \int_{0}^{\frac{3 \pi}{2}} \sin^3 x \cos^2 x dx\) i tried solving it but the answer did not match the choices given..i think i did something wrong

OpenStudy (lgbasallote):

lemme fix the latex

OpenStudy (blockcolder):

Can you show us your solution?

OpenStudy (lgbasallote):

okay..gimme some time

OpenStudy (mimi_x3):

Write it as: \[\int\limits sinxsin^{2}xcos^{2}x =>\int\limits sinx(1-\cos^{2}x)(\cos^{2}x) \] Might be easier now..it's a substitution..

OpenStudy (lgbasallote):

\(\int (\sin ^2x)(\cos ^2 x) \sin x dx\) \(\int (1 - \cos^2 x)(\cos ^2 x) \sin x dx\) \(\int (\cos ^2 x - \cos ^4 x)\sin x dx\) let u = \(\cos x\) du = \(-\sin x dx\) \(-\int (u^2 - u^4) du\) \(\LARGE \frac{-\cos ^3 x}{3} + \frac{\cos ^5 x}{5}\)

OpenStudy (mimi_x3):

looks right..

OpenStudy (lgbasallote):

hmm seems i got it...guess i just needed to put it in latex :P

OpenStudy (mimi_x3):

lol, yeah then sub in the limit

OpenStudy (lgbasallote):

\(\huge \frac{-[\cos (\frac{3\pi}{2})]^3}{3} + \frac{[\cos (\frac{3\pi}{2})]^5}{5}\)

OpenStudy (blockcolder):

You forgot 0. cos(0)=1. :D

OpenStudy (lgbasallote):

oh..oh yeah..silly me...no wonder i couldnt get this

OpenStudy (lgbasallote):

uhmmm hypothetical question...what's cos(3pi/2)? this is kinda ironic...considering i did a tutorial on it not so long ago :P

OpenStudy (mimi_x3):

0 lol you are able to use a calculator :P

OpenStudy (lgbasallote):

hmph

OpenStudy (mimi_x3):

why hmph-ing? like me :p

OpenStudy (lgbasallote):

coz you wouldnt teach me how to do it :P

OpenStudy (mimi_x3):

im your teacher now huh :P

OpenStudy (lgbasallote):

you said youll guide me :P it's just the last step..just cos (3pi/2)

OpenStudy (mimi_x3):

need anymore help?

OpenStudy (lgbasallote):

i dont know...im just kind of thinking a new topic for my next tutoriall..you?

OpenStudy (mimi_x3):

i gotta go; the next tutorial can be on the trig inverses idk